OpenStudy (happpily):

what are the solutions 1/2x^2+2x+3=0 ?

OpenStudy (danjs):

multiply everything by 2 to get rid of that fraction

OpenStudy (danjs):

\[x^2 + 4x + 6 = 0\]

OpenStudy (danjs):

(x + )(x + ) = 0

OpenStudy (danjs):

what are the factors of 6?

OpenStudy (danjs):

you need two numbers that when multiplied together give you 6, but when added together give you 4

geerky42 (geerky42):

And there are no such integers.

OpenStudy (danjs):

or use the quadratic formula

geerky42 (geerky42):

^ I recommend that

OpenStudy (happpily):

wait just solved it i got x = -2+i squareroot2, x = -2- i squareroot 2 is that right?

geerky42 (geerky42):

Looks good.

OpenStudy (happpily):

yeah i used the quadratic formula :D!

OpenStudy (danjs):

\[x = \frac{ -b +- \sqrt{b^2 - 4*a*c} }{ 2a }\] a=1 b=2 c=6

OpenStudy (danjs):

cool

OpenStudy (happpily):

awesome thanks guys :D

OpenStudy (danjs):

i got something different

geerky42 (geerky42):

Try simplify your answer

OpenStudy (danjs):

x = -2 + sqrt(2) i x = -2 - sqrt(2) i

geerky42 (geerky42):

Yeah, this is what @Happpily got.

OpenStudy (danjs):

oh, lol i didnt read the "squareroot" thing, beg my pard